# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v1

@@ 3.1 Univariate generalized extreme value (UGEV)
As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most safety-critical near-miss events in both V–V and V–I interactions. To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s). Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001). The resulting sequence of BM serves as the input for parameter estimation. Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$. To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
@@ 3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model. Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level. This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability. In this specification, covariates are partitioned into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups. For a dataset divided into $K$ groups, with observation $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group. This enables the capture of localized relationships while preserving overall model coherence. The model is expressed within a three-layer hierarchical Bayesian framework. The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution. Process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence. Prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
@@
In the above equations $X_{k,j}$ defined as group-level fixed-effect covariate $j$, $W_{k,i,j}$is sample-specific random-effect covariate $j$. $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters. $\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients. $\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$. Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
@@ 3.2 COR estimation
Building on the HBSGRP–UGEV estimates, the framework compute COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block. Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, enabling a holistic assessment of potential COR under mixed traffic conditions.
Let $i$ denote block $i$ within group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments). The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone. The probability of a crash-level event in block $i$ of group $k$ is:
@@
Assuming a homogeneous Poisson process for exceedances, the expected COR for group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$. Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$. The expected COR for group $k$ is:
@@ 4 Data description
The effectiveness of real-time COR estimation depends critically on the fidelity and coverage of trajectory data. Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013). Recent advances in AV sensing offer a promising alternative. AV datasets offer high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers. This richness allows for realistic near-miss detection in complex traffic environments. Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation. For example, studies using the Waymo Open Dataset have largely focused on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
@@
To address these challenges, this study utilizes the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared with Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions such as intersection near-misses, lane changes, and pedestrian crossings—conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
Table 1: Open-source AV datasets
Dataset Sensora Tasksb Hzc HD Map Scope / Notes
Argoverse-2 (Wilson et al. 2023) C, L D, T, S, M 10 Yes 250k scenarios across 113 segments; 11s each; 10 classes; 6 cities
nuScenes (Caesar et al. 2020) C, L, R D, T, S, M 2 No 40k scenarios across  1k segments; 23 classes; 2 cities
Waymo (Ettinger et al. 2021) C, L D, T, M 10 Yes 390k scenarios over  1950 segments; 12.6M trajectories; 6 cities
Lyft Level 5 (Kesten et al. 2019) C, L, R M 10 Yes 170k AV instances (25s) in one city; 10 classes
KITTI (Geiger et al. 2013) L D, T, M 10 No 15k scenes; classic full-stack AV benchmark
DAIR-V2X (Yu et al. 2022) L D Yes 71254 image frames for V2X scenes
@@
OPV2V (Xu et al. 2022) C, L D, T, M 10 No 33k V2V simulation samples; 18k V2X frames (CARLA)
DeepAccident (Wang et al. 2024) C, L, R D, T, S, M 10 Yes 285k simulation samples incl. 57k V2X scenarios
Shift (Sun et al., 2022) C, L S, M 10 No 2.5M frames; 4850 segments; 8 cities
@@
    Sensor: Camera (C), LiDAR (L), Radar (R).
@@
    Tasks: Detection (D), Tracking (T), Segmentation (S), Motion Forecasting (M).
  • c
    Hz: Sampling frequency; “–” not reported.

### v2

@@ 3.1 Univariate generalized extreme value (UGEV)
As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most critical near-miss events in both V–V and V–I interactions. To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s). Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001). The resulting sequence of BM serves as the input for parameter estimation. Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$. To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
@@ 3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model. Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level. This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability. In this specification, covariates are divided into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups. For a dataset divided into $K$ groups, with observations $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group. This enables the capture of localized relationships while preserving overall model coherence. The model is expressed within a three-layer hierarchical Bayesian framework. The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution. The process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence. The prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
@@
In the above equations, $X_{k,j}$ defined as a group-level fixed-effect covariate $j$, $W_{k,i,j}$is a sample-specific random-effect covariate $j$. $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters. $\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients. $\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$. Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
@@ 3.2 COR estimation
Building on the HBSGRP–UGEV estimates, the framework computes COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block. Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, allowing for a holistic assessment of potential COR under mixed traffic conditions.
Let $k,i$ denote the block $i$ within the group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments). The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone. The probability of a crash-level event in block $i$ of group $k$ is:
@@
Assuming a homogeneous Poisson process for exceedances, the expected COR for the group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$. Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$. The expected COR for group $k$ is:
@@ 4 Data description
The effectiveness of real-time COR estimation is critically dependent on the fidelity and coverage of trajectory data. Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013). Recent advances in AV sensing offer a promising alternative. AV datasets provide high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers. This richness allows for realistic near-miss detection in complex traffic environments. Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation. For example, studies using the Waymo Open Dataset have focused mainly on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
@@
To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
Table 1: Representative open-source AV datasets with applications in safety and crash-risk analysis
Dataset Trajectory coveragea Safety relevanceb Duration / Frequency HD map Notes
Argoverse-2 250k scenarios; 113 segments; 6 cities Near-miss analysis, surrogate safety measures 11s per scenario, 10 Hz Yes Rich lane-level geometry; 10 object classes (Wilson et al. 2023)
nuScenes 40k scenes; $\sim$1000 segments; 2 cities Multi-agent interactions, trajectory risk indicators 20s per scene, 2 Hz No 23 object classes; widely used urban dataset (Caesar et al. 2020)
Waymo 390k scenarios; $\sim$1950 segments; 6 cities Motion forecasting, crash-prone interaction detection Up to 20s, 10 Hz Yes 12.6M trajectories across diverse geometries (Ettinger et al. 2021)
Lyft Level 5 170k instances; 1 city Vehicle interactions; trajectory-based safety studies 25s per instance, 10 Hz Yes Limited geographic diversity; 10 object classes (Kesten et al. 2019)
KITTI 15k scenes; 1 city Early surrogate measures, safety benchmarks 1–2 min sequences, 10 Hz No Classic full-stack dataset, still widely cited (Geiger et al. 2013)
DAIR-V2X 71k frames Infrastructure-based safety monitoring Yes First large-scale real-world V2X dataset (Yu et al. 2022)
@@
OPV2V 33k samples; CARLA simulation Cooperative perception, proactive safety design 10 Hz No 18k V2X frames, sim environment (Xu et al. 2022)
DeepAccident 285k samples; sim dataset Rare crash/near-crash events, proactive safety 10 Hz Yes Includes 57k V2X scenarios (Wang et al. 2024)
Shift 2.5M frames; 4850 segments; 8 cities Long-term safety and mobility analysis 10 Hz No Corridor-scale benchmarking dataset (Sun et al., 2022)
@@
    Trajectory coverage: Number of driving scenarios and geographic spread.
@@
    Safety relevance: Potential application for crash risk, near-miss, or surrogate safety studies.